OpenStudy (well):

X,Y ?! (3+2i)x - 5(y+3i)+2i = y(3-7i) - x(1-3i)-3i. Y should be 2 and X = 4. But i dont know steps.

OpenStudy (asnaseer):

Expand each side and then equate the real and imaginary components. This should give you two equations in two unknowns (x and y), which you can solve to find x and y.

OpenStudy (asnaseer):

e.g. if you have: a + ib = c + id Then this implies: a = c b = d

OpenStudy (well):

So, I have 3x - 5y = 3y - 1x 2x - 15 + 2 = -7y +3x -3 Right ?

OpenStudy (asnaseer):

Yes - that is correct so far

OpenStudy (well):

Can i modify it on ? : 4x - 8y = 0 -1x + 7y = 10

OpenStudy (asnaseer):

That is still correct

OpenStudy (asnaseer):

note that you can divide the 1st equation by 4

OpenStudy (well):

x - 2y = 0 -x +7y = 10

OpenStudy (asnaseer):

yup - now just solve this pair to find x and y

OpenStudy (well):

5y = 10 y = 2 Now i just take Y into 1st equation ?

OpenStudy (asnaseer):

perfect!

OpenStudy (well):

Thanks a lot. I dont know where did mistakes when i try it solve alone. Perfect page ! Thanks

OpenStudy (asnaseer):

yw - and welcome to OpenStudy! :)

OpenStudy (well):

now just 19 examples to do , on complex numbers :D

OpenStudy (asnaseer):

:)

OpenStudy (well):

I hope that I can handle it myself :)

OpenStudy (asnaseer):

Just so you are aware, the standard workflow here is as follows: 1. You post a question. 2. One or more people come to your question to help you out. 3. You select the one that best answered your question by clicking on the blue Best Response button next to any one of their replies. 4. You close the question by clicking on the Close button then the whole process starts again.

OpenStudy (well):

Ok, thanks for your time.

OpenStudy (asnaseer):

again - yw :)